OpenStudy (anonymous):

reduce to slope-intercept form : 3x-4y=-24

OpenStudy (anonymous):

subtract the 3x to the other side the divide the equation by -4

OpenStudy (anonymous):

it should look like this y= 3/4x +6

OpenStudy (anonymous):

can u write the whole solution so that i can check mine please

Miracrown (miracrown):

Slope-intercept form is when y is all by itself on one side of the equation. The number in front of the x is the slope, with the number being added or subtracted from the x is the y-intercept. Hence the name, slope-intercept form. So, try and get y all by itself on the left. left side of the equation.

OpenStudy (anonymous):

3x-4y+-24 subtract 3x -4y=-3x-24 divide by -4 y=3/4+6 is that how you want it

OpenStudy (anonymous):

then how do i graph it ?

OpenStudy (anonymous):

so you go down -6 on the y-axis then look at your slope and go from there. so you would go up 3 on the y-axis then over 4 on the x-axis

OpenStudy (anonymous):

in your solution , what do u mean divide by 4

OpenStudy (anonymous):

 is it this ?